Session: Feb 1998 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Daniele Dallacasa at Sun Feb 22 14:55:00 MET 1998) Success
Very dense fog till 9:00 ut, then light clouds. No problem during observation.

Onsala (A. Polatidis at Sun Feb 22 19:24:41 MET 1998) Success
No known problems

Torun (K. Borkowski at Mon Feb 23 10:21:19 MET 1998) Success
No known problems

Westerbork (Tony Foley at Mon Feb 23 11:03:42 MET 1998) Success
Looks fine, but total power levels were adjusted to get good values on CygA, and so are low on 1823+568 and 2005+403.

Jodrell (Alastair Gunn at Mon Feb 23 11:18:33 MET 1998) Minor failures
Extreme RFI in VC 9, receiver often saturated. Some attempt made to alter attenuation levels between sources. Patchy cloud, some rain. No other known problems.

Noto (C.Trigilio at Wed Feb 25 14:30:58 MET 1998) Minor failures
BBC08 not working fine; unlocked for almost all the experiment. Lost from 11:13 to 11:46 for cable wrap problems.

Urumqi (Liu Xiang at Fri Feb 27 05:01:08 MET 1998) Success

Calibration (jean\-francois Desmurs at Mon Mar 2 18:49:57 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb98 directory under the name ea11a.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Thu Mar 5 16:01:54 MET 1998) Minor failures
1823+568 at 0745 lost. rain. total power level offscale.